# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v2

@@ 1.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths, over 615 injuries, and widespread destruction throughout the city. A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with the improper storage of hazardous materials.
@@ 1.3 Volcano Eruption in DR Congo and Rwanda, 2021
On May 22, 2021, Mount Nyiragongo in the Democratic Republic of the Congo erupted, causing widespread devastation. The eruption resulted in 32 deaths, the destruction of 1,000 homes, and the displacement of thousands as lava flows threatened the city of Goma. Nearly 400,000 people were evacuated due to the risk of further volcanic activity, including potential magma flow beneath Goma and nearby Lake Kivu. Despite continued seismic activity, life in Goma largely returned to normal by August 2021, although plans to relocate parts of the city remain under consideration due to the ongoing threat from the volcano.
@@ 1.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths, more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings, including homes, schools, and hospitals, were damaged or destroyed. As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
@@ 1.5 Volcano Eruption in La Palma, 2021
On September 19, 2021, the Cumbre Vieja volcano on La Palma, part of Spain’s Canary Islands, erupted following several days of seismic activity. The eruption primarily impacted the island’s western side, covering over 1,000 hectares with lava and destroying more than 3,000 buildings, including the towns of Todoque and La Laguna. The lava flow, measuring about 3.5 kilometers wide and 6.2 kilometers long, reached the sea, cutting off the coastal highway and forming a new peninsula with extensive lava tubes. Although the timely evacuation of around 7,000 people prevented major casualties, one person died from inhaling toxic gases. Economic losses exceeded 800 million Euros, and thousands of residents were displaced, with significant damage to arable land and livelihoods.
@@ 1.6 Wildfire in Colorado, 2021
The Marshall Fire, which started on December 30, 2021, in Boulder County, Colorado, became the most destructive wildfire in the state history in terms of destroyed buildings. Fueled by dry grass from an unusually warm and dry season, and winds up to 185 km/h, the fire killed two people and injured eight. It destroyed 1,084 structures, including homes, a hotel, and a shopping center, causing over US$2 billion in damage. More than 37,500 residents evacuated, and significant damage to public drinking water systems occurred.
@@ 1.7 Armed Conflict in Ukraine, 2022
In February 2022, the conflict in Ukraine escalated significantly when Russian forces launched a full-scale invasion. The invasion has affected many parts of Ukraine, especially the eastern regions and the major cities, causing widespread destruction to homes, infrastructure, and industrial facilities. The conflict has resulted in thousands of civilian and military casualties and has caused the displacement of millions of people. Out of a population of 41 million, about 8 million Ukrainians were internally displaced, while more than 8.2 million had fled the country by April 2023, creating Europe’s largest refugee crisis since World War II. The economic losses are substantial, with extensive damage to infrastructure, industrial output, and the broader economy.
@@ 1.8 Earthquake in Turkey, 2023
On February 6, 2023, a magnitude 7.8 earthquake struck southeastern Turkey near Gaziantep, followed by a magnitude 7.7 aftershock. The disaster, the most powerful earthquake in Turkey since 1939, caused widespread destruction across approximately 350,000 , affecting 14 million people and displacing 1.5 million. The death toll reached 53,537 in Turkey and 5,951 in Syria, with 107,213 injuries, making it one of the deadliest earthquakes in modern history. Economic losses were estimated at USD 148.8 billion in Turkey and USD 14.8 billion in Syria, with over 518,000 houses and 345,000 apartments destroyed. The earthquake caused severe damage to infrastructure, agriculture, and essential services, further worsening the region’s economic challenges. International aid was mobilized to support the affected populations.
@@ 1.9 Cyclone in Myanmar, 2023
In May 2023, Cyclone Mocha, a Category 4 hurricane, struck Myanmar, causing widespread devastation in the country’s coastal regions, particularly in Rakhine State. The storm resulted in at least 460 deaths, with over 700 people injured, mostly among the Rohingya refugees. The cyclone destroyed over 183,000 houses, 1,770 religious buildings, and 1,397 schools, and caused US$2.24 billion in damages, equivalent to 3.4% of Myanmar’s GDP. Severe flooding affected 895 of land, displacing over 1.2 million people. In Rakhine, towns like Sittwe were heavily flooded, with a storm surge of up to 3.5 meters, while nearly 90% of homes in Kyauktaw were destroyed. The cyclone severely affected communication networks and infrastructure, exacerbating the region’s vulnerability to future natural disasters.
@@ 1.10 Wildfire in Hawaii, 2023
In August 2023, a series of wildfires broke out on the island of Maui, Hawaii, causing widespread destruction and significant impacts on the local population and environment. The fires, fueled by dry conditions and strong winds, primarily affected the town of Lahaina, where at least 102 people were killed and two remain missing. Over 2,200 buildings were destroyed, including many historic landmarks, resulting in estimated damages of US$5.5 billion. The fires prompted evacuations and led to the displacement of thousands of residents, with significant economic losses in the tourism and agriculture sectors.
@@ 1.11 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province near Marrakesh, causing widespread devastation. The quake resulted in at least 2,960 deaths and 5,674 injuries, affecting over 2.8 million people, including 100,000 children. The earthquake damaged or destroyed 40,759 houses and 2,930 villages, with 19,095 additional houses collapsing. Significant destruction occurred in rural areas of the Atlas Mountains. Economic losses could reach up to 9% of Morocco’s GDP. Historic sites in Marrakesh, including parts of the Medina and several mosques, were severely damaged.
@@ 1.12 Flood and Storm in Libya, 2023
In September 2023, northeastern Libya was hit by Storm Daniel, causing catastrophic flooding, particularly in the city of Derna. The storm led to the collapse of two dams, releasing 30 million cubic meters of water and partially destroying the city. Casualty estimates range from 5,900 to 20,000, making it the second deadliest dam failure in history. The flooding affected seven out of 10 districts in Derna and left more than 40,000 people displaced. More than 2,200 buildings were flooded and the collapse of four bridges split the city into two. The economic losses were substantial, with severe damage to infrastructure, homes, and the agricultural sector.
@@ 1.13 Hurricane in Mexico, 2023
In September 2023, Hurricane Norma, a Category 4 hurricane, struck the western coast of Mexico, severely affecting Sinaloa and Baja California Sur. The storm caused over 100 deaths and displaced thousands, with significant damage to homes, infrastructure, and agriculture. This was followed by Hurricane Otis in October, which made landfall near Acapulco as a Category 5 hurricane. Otis was the strongest Pacific hurricane to hit Mexico, causing at least 52 deaths and leaving 32 missing. The storm caused unprecedented destruction, with more than 51,864 homes destroyed and damages estimated at US$12–16 billion, surpassing Hurricane Wilma as the costliest Mexican hurricane.
@@ 1.14 Earthquake in Noto, 2024
On January 1, 2024, a magnitude 7.5 earthquake struck the Noto Peninsula in Ishikawa Prefecture, Japan, reaching a maximum JMA seismic intensity of Shindo 7. The earthquake caused widespread destruction, particularly in the towns of Suzu, Wajima, Noto, and Anamizu, and triggered a 6.58-meter tsunami along the Sea of Japan coast. The disaster resulted in 339 deaths and over 1,300 injuries, with 156,526 structures damaged across nine prefectures. It was the deadliest earthquake in Japan since the 2011 Tōhoku disaster. As of February 20, 12,929 people remained in 521 evacuation centers. The event prompted Japan’s first major tsunami warning since 2011. Economic losses are estimated to be substantial, with significant impacts on infrastructure and communities across the affected regions.
@@ 2.1 Construction of the dataset
Figure 3 shows the flowchart of developing BRIGHT. The optical EO data in our dataset are mainly from Maxar’s Open Data program333https://www.maxar.com/open-data, while the SAR EO data are from Capella Space444https://www.capellaspace.com/earth-observation/gallery and Umbra Space555https://umbra.space/open-data/. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if there is suitable data in the data provider’s inventory. Otherwise, we chose Stripmap. The optical EO data consist of red, blue, and green bands, while the SAR EO data are amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data had been terrain-corrected, we used the data provider’s recommended method of pre-processing the amplitude data and finally converted the data to an 8-bit data format. Although both optical and SAR images are geocoded, there are still pixel offsets between them. Therefore, multiple EO experts manually aligned the paired optical and SAR data and cross-checked their results to ensure the precise registration between the two types of EO data. Figure 11 shows the selected control points on three disaster scenes.
| Class | Definition |
@@
The labels in BRIGHT consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, and all labels underwent independent visual inspections to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service666https://emergency.copernicus.eu, the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products777https://unosat.org/products, and The Federal Emergency Management Agency (FEMA)888https://www.fema.gov. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. To further improve damage information reliability, we performed a secondary validation and correction of these annotations. EO experts carefully reviewed the damage points using multi-temporal imagery available on platforms such as Google Earth and manually refined the labels when inconsistencies or misclassifications were observed. Since the number of damage levels varied across different events, we refered to the framework in FEMA’s Damage Assessment Operations Manual, EMS-98, the BDD dataset (Adriano et al., 2021), and the xBD dataset (Gupta et al., 2019) to standardize them into three categories: intact (1), damaged (2), and destroyed (3). Table 3 summarizes these definitions. The damage annotations were provided as vector point files. By overlaying these points with the building polygons and assigning corresponding damage attributes, we generated the final building damage labels. To prevent geographic misallocations due to possible coordinate offsets, we unified the coordinate systems of the points and polygons, with a visual inspection performed prior to the final allocation.
Figure 4 presents thumbnails of selected local areas from the 14 disaster events included in the BRIGHT dataset.
@@ 2.2 Statistics for BRIGHT
The key statistics of BRIGHT are illustrated in Figure 5. Figure 5a shows the pixel value distribution for optical and SAR images from one human-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values, highlighting BRIGHT’s geographical diversity and making it a robust dataset for studying building damage assessment in diverse environments. To ensure trained deep learning (DL) models can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 5b shows that BRIGHT covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
@@
We also evaluated the registration accuracy between optical and SAR EO data in BRIGHT. Due to the absence of real ground truth, we introduce a proxy method to estimate registration errors, leveraging existing multimodal image descriptors. This approach is detailed in Appendix B. Table 4 reports the mean registration errors, measured as the root mean square error (RMSE) of pixel displacements, obtained using three representative multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The overall average RMSE is approximately 1.024 pixels, with a lower error of 1.006 pixels specifically within building regions. Accurate registration results ensure spatial consistency across modalities, providing a solid foundation for accurate building damage assessment in BRIGHT.

### v3

@@ 1.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths, over 615 injuries, and widespread destruction throughout the city (OCHA, 2021a). A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with the improper storage of hazardous materials.
@@ 1.3 Volcano Eruption in DR Congo and Rwanda, 2021
On May 22, 2021, Mount Nyiragongo in the Democratic Republic of the Congo erupted, causing widespread devastation. The eruption resulted in 32 deaths, the destruction of 1,000 homes, and the displacement of thousands as lava flows threatened the city of Goma (IFRC, 2021). Nearly 400,000 people were evacuated due to the risk of further volcanic activity, including potential magma flow beneath Goma and nearby Lake Kivu. Despite continued seismic activity, life in Goma largely returned to normal by August 2021, although plans to relocate parts of the city remain under consideration due to the ongoing threat from the volcano.
@@ 1.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths, more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings, including homes, schools, and hospitals, were damaged or destroyed (OCHA, 2021b). As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
@@ 1.5 Volcano Eruption in La Palma, 2021
On September 19, 2021, the Cumbre Vieja volcano on La Palma, part of Spain’s Canary Islands, erupted following several days of seismic activity. The eruption primarily impacted the island’s western side, covering over 1,000 hectares with lava and destroying more than 3,000 buildings, including the towns of Todoque and La Laguna. The lava flow, measuring about 3.5 kilometers wide and 6.2 kilometers long, reached the sea, cutting off the coastal highway and forming a new peninsula with extensive lava tubes. Although the timely evacuation of around 8,000 people prevented major casualties, one person died from inhaling toxic gases (Troll et al., 2024). Economic losses exceeded 800 million Euros, and thousands of residents were displaced, with significant damage to arable land and livelihoods.
@@ 1.6 Wildfire in Colorado, 2021
The Marshall Fire, which started on December 30, 2021, in Boulder County, Colorado, became the most destructive wildfire in the state history in terms of destroyed buildings. Fueled by dry grass from an unusually warm and dry season, and winds up to 185 km/h, the fire killed two people and injured eight. It destroyed 1,084 structures, including homes, a hotel, and a shopping center, causing over US$2 billion in damage (Forrister et al., 2024). More than 37,500 residents evacuated, and significant damage to public drinking water systems occurred (Forrister et al., 2024).
@@ 1.7 Armed Conflict in Ukraine, 2022
In February 2022, Russian forces launched a full-scale invasion of Ukraine, resulting in widespread destruction and displacement. By November 2024, total damages to Ukraine’s infrastructure had reached $170 billion. More than 236,000 residential buildings were damaged or destroyed, including 209,000 private houses, 27,000 apartment buildings, and 600 dormitories. Over 4,000 educational institutions and 1,554 medical facilities were also affected, alongside extensive damage to transport, energy, and telecommunications infrastructure. The conflict has inflicted more than 40,000 civilian casualties, displaced four million people internally, and forced 6.8 million refugees to flee. As of late 2024, approximately 14.6 million Ukrainians require humanitarian assistance (Andrienko et al., 2025).
@@ 1.8 Earthquake in Turkey, 2023
On February 6, 2023, a magnitude 7.8 earthquake struck southeastern Turkey near Gaziantep, followed by a magnitude 7.7 aftershock. The disaster, the most powerful earthquake in Turkey since 1939, caused widespread destruction across approximately 350,000 , affecting 14 million people and displacing 1.5 million. The death toll reached 53,537, with 107,213 injuries (STL, 2024), making it one of the deadliest earthquakes in modern history. Economic losses were estimated at USD 148.8 billion, with over 518,000 houses and 345,000 apartments destroyed (Government of Türkiye, 2023). The earthquake caused severe damage to infrastructure, agriculture, and essential services, further worsening the region’s economic challenges. International aid was mobilized to support the affected populations.
@@ 1.9 Cyclone in Myanmar, 2023
In May 2023, Cyclone Mocha, a Category 5 hurricane, struck Myanmar, causing widespread devastation in the country’s coastal regions, particularly in Rakhine State. According to official estimates, at least 148 people were killed and 132 injured, although other sources suggest higher figures (GFDRR and World Bank, 2023). The cyclone affected around 1.2 million people in Rakhine alone, with over 200,000 buildings reportedly damaged or destroyed, making it one of the most destructive cyclones in the region in the past 15 years. The direct economic damage was estimated at US$2.24 billion in damages, equivalent to 3.4% of Myanmar’s GDP (GFDRR and World Bank, 2023).
@@ 1.10 Wildfire in Hawaii, 2023
In August 2023, a series of wildfires broke out on the island of Maui, Hawaii, causing widespread destruction and significant impacts on the local population and environment. The fires, fueled by dry conditions and strong winds, primarily affected the town of Lahaina, where at least 102 people were killed and two remain missing (Hedayati et al., 2024). Over 2,200 buildings were destroyed, including many historic landmarks, resulting in estimated damages of US$5.5 billion (Jones et al., 2024; NCEI, 2025). The fires prompted evacuations and led to the displacement of thousands of residents, with significant economic losses in the tourism and agriculture sectors.
@@ 1.11 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province near Marrakesh, causing widespread devastation. The earthquake killed nearly 3,000 people, injured more than 5,500 and displaced more than half a million (OCHA, 2023). The earthquake damaged or destroyed nearly 60,000 houses, with the heaviest losses reported in rural communities of the Atlas Mountains (International Federation of Red Cross and Red Crescent Societies, 2024). Overall damage is estimated at about USD 7 billion (National Geophysical Data Center / World Data Service, 2025), while direct economic losses of roughly USD 30 million amount to approximately of Morocco’s GDP. Cultural heritage in Marrakesh was also hard-hit: sections of the UNESCO-listed Medina and several historic mosques sustained severe damage.
@@ 1.12 Flood and Storm in Libya, 2023
In September 2023, Storm Daniel brought catastrophic flooding to northeastern Libya, particularly in Derna, when two dams collapsed and released an estimated 30 million cubic meters of water. At least 5,923 people were killed, though local officials warned fatalities could reach 18,000–20,000 (USAID Bureau for Humanitarian Assistance, 2024). Post–event analyses revealed that approximately of Derna’s housing stock was destroyed and a further sustained damage. Across the wider coastal belt—from Benghazi, through Jabal Al Akhdar and Al Marj, to Derna—an estimated 18 838 dwellings were damaged or obliterated (Normand and Heggy, 2024).. The disaster, considered the second deadliest dam failure in history, devastated the city’s infrastructure: four bridges collapsed, and entire neighborhoods were washed out to sea. Long-standing neglect of dam maintenance, compounded by Libya’s political turmoil, contributed significantly to the scale of the tragedy.
@@ 1.13 Hurricane in Mexico, 2023
In September 2023, Hurricane Norma, a Category 4 hurricane, struck the western coast of Mexico, severely affecting Sinaloa and Baja California Sur. This was followed by Hurricane Otis in October, which made landfall near Acapulco as a Category 5 hurricane. Otis was the strongest Pacific hurricane to hit Mexico, causing at least 52 deaths and leaving 32 missing. The storm caused unprecedented destruction, with more than 51,864 homes destroyed and damages estimated at US$12–16 billion, surpassing Hurricane Wilma as the costliest Mexican hurricane (Reinhart and Reinhart, 2024).
@@ 1.14 Earthquake in Noto, 2024
On January 1, 2024, a magnitude 7.5 earthquake struck the Noto Peninsula in Ishikawa Prefecture, Japan, reaching a maximum JMA seismic intensity of Shindo 7. The earthquake caused widespread destruction, particularly in the towns of Suzu, Wajima, Noto, and Anamizu, and triggered a 7.45-meter tsunami along the Sea of Japan coast. The disaster resulted in 572 deaths and over 1,300 injuries, with 193,529 structures damaged across nine prefectures (Fire and Disaster Management Agency, 2025). It was the deadliest earthquake in Japan since the 2011 Tōhoku disaster. As of February 20, 12,929 people remained in 521 evacuation centers (Japanese Red Cross Society, 2024). The event prompted Japan’s first major tsunami warning since 2011. Total damage is estimated 16.9 trillion yen, reflecting substantial economic losses with significant impacts on infrastructure and communities across the affected regions (Kaneko, 2024).
@@ 2.1 Construction of the dataset
Figure 3 shows the flowchart of developing Bright. The optical EO data in our dataset are mainly from Maxar’s Open Data program333https://www.maxar.com/open-data, while the SAR EO data are from Capella Space444https://www.capellaspace.com/earth-observation/gallery and Umbra555https://umbra.space/open-data/. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if there is suitable data in the data provider’s inventory. Otherwise, we chose Stripmap. The optical EO data consists of red, blue, and green bands, while the SAR EO data are amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data had been terrain-corrected, we used the data provider’s recommended method of pre-processing the amplitude data and finally converted the data to an 8-bit data format. Although both optical and SAR images are geocoded, there are still pixel offsets between them. Therefore, multiple EO experts manually aligned the paired optical and SAR data and cross-checked their results to ensure the precise registration between the two types of EO data. Figure 11 shows the selected control points on three disaster scenes.
| Category | Definition |
@@
The labels in Bright consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, and all labels underwent independent visual inspections of EO experts to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service666https://emergency.copernicus.eu, the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products777https://unosat.org/products, and The Federal Emergency Management Agency (FEMA)888https://www.fema.gov. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. To further improve damage information reliability, we performed a secondary validation and correction of these annotations. EO experts carefully reviewed the damage points using multi-temporal imagery available on platforms such as Google Earth and manually refined the labels when inconsistencies or misclassifications were observed. Since the number of damage levels varied across different events, we refered to the framework in FEMA’s Damage Assessment Operations Manual, EMS-98, the BDD dataset (Adriano et al., 2021), and the xBD dataset (Gupta et al., 2019) to standardize them into three categories: intact (1), damaged (2), and destroyed (3). Table 3 summarizes these definitions. The damage annotations were provided as vector point files. By overlaying these points with the building polygons and assigning corresponding damage attributes, we generated the final building damage labels. To prevent geographic misallocations due to possible coordinate offsets, we unified the coordinate systems of the points and polygons, with a visual inspection performed prior to the final allocation.
Figure 4 presents thumbnails of selected local areas from the 14 disaster events.
@@ 2.2 Statistics for Bright
The key statistics of Bright are illustrated in Figure 5. Figure 5a shows the pixel value distribution for optical and SAR images from one human-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values, highlighting Bright’s geographical diversity and making it a robust dataset for studying building damage assessment in diverse environments. To ensure trained models can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 5b shows that Bright covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
@@
We also evaluated the registration accuracy between optical and SAR EO data in Bright. Due to the absence of real ground truth, we introduce a proxy method to estimate registration errors, leveraging existing multimodal image descriptors. This approach is detailed in Appendix B. Table 4 reports the mean registration errors, measured as the root mean square error (RMSE) of pixel displacements, obtained using three representative multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The overall average RMSE is approximately 1.024 pixels, with a lower error of 1.006 pixels specifically within building regions. Accurate registration results ensure spatial consistency across modalities, providing a solid foundation for accurate building damage assessment.